TTFDUMP(1) TTFDUMP(1) NAME ttfdump−DumpingthecontentsofaTrueTypeFontfile SYNOPSIS ttfdump[-h][-ttablename][-gglyphnumber][-ccollection][-i]ttfile[-odumpfile] DESCRIPTION ttfdumpdumpsthecontentsofaTrueTypefontfileinASCIIform.ATrueTypefontfileisconsistofvarioustables.Thosetablearedividedintotwocategories-RequiredTablesandOptional Tables.MulitipleTrueTypefontsandbemergedtoaTrueTypeCollection. Options -hPrintahelpmessageonstandardoutputandexit. -ttablename Ifthisoptionisspecifiedthendumpthetable"tablename",otherwisedumpeverytables inttfile.Tablenamecanbeoneofthefollowing: RequiredTables cmapcharactertoglhphmappingtable glyfglyphdata headfontheader hheahorizontalheader hmtxhorizontalmetrics locaindextolocation maxpmaximumprofile namenamingtable postPostScriptinformation OS/2OS/2andwindowsspecificmetrics OptionalTables cvtcontrolvaluetable fpgmfontprogram;notimplementedyet. gaspgrid-fittingandscanconversionprocedure(grayscale) hdmxhorizontaldevicemetrics kernkerningtable LTSHLinearthresholdtable prepCVTprogram;notimplementedyet PCLTPCL5table VDMX VerticalDeviceMetricstable vheaverticalmetricsheader vmtxverticalmetrics -gglyphnumber Ifthisoptionisspecifiedthendumptheglyphwithindexglyphnumber,otherwisedump allglyphsinttfile. -ccollection SelectaTrueTypefontinaTrueTypeCollection. ILXProject 18Feb1997 1 TTFDUMP(1) TTFDUMP(1) -ittfileTrueTypefontfiletobedumped. -odumpfile Outputofthedump.Ifthisoptionisnotspecified,dumptostdout. EXAMPLE TodumpallinformationintheTrueTypefontfiletimes.ttf: ttfdumptimes.ttf Todumptherequiredtablecmapinthefile: ttfdump-tcmaptimes.ttf Todumpasingleglyphwithindex50: ttfdump-g50times.ttf SEEALSO TrueType1.0FontFiles,TechnicalSpecificationRevision1.66November1995 NOTE Instructiondisassemblerinnotimplemented. AOTHOR Li-DaLho,ollie@ms1.hinet.net ILXProject 18Feb1997 2